OpenStudy (anonymous):

find the slope intercept equation of the line that has given characteristics slope is 5 and y intercept =

OpenStudy (saifoo.khan):

y-intercept ???

OpenStudy (saifoo.khan):

its missing i guess.

OpenStudy (anonymous):

slope 5 and what is the y intercept ? y=

OpenStudy (saifoo.khan):

there is still something missing,

OpenStudy (saifoo.khan):

there might b the co-ordinated have been given.

OpenStudy (anonymous):

it says to find the slope intercept of the line that has given characteristics slope 5 and y intercept (0,9)

OpenStudy (saifoo.khan):

Yeah. not it is solvable! ^.^

OpenStudy (anonymous):

the slope intercept equation is y =

OpenStudy (saifoo.khan):

y = mx + c 9 = 0 + c c = 9 therefore: y = 5x + 9

OpenStudy (anonymous):

thank you do you think you can help me with a couple more please ?

OpenStudy (saifoo.khan):

Sure. :)

OpenStudy (anonymous):

they are a bit harder :(

OpenStudy (saifoo.khan):

Lets Try.

OpenStudy (anonymous):

k the function p(d) =1+d/33 gives the pressure in atmospheres (atm) at a depth d in the sea (d) is in feet note that p(o) =1 (atm) p 33=2 find the pressure at 140 feet is blank (atm)

OpenStudy (anonymous):

see its hard

OpenStudy (saifoo.khan):

Reading....

OpenStudy (saifoo.khan):

is it p(d) = 1 +(d/33) right?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

chk this sentence too: note that p(o) =1 (atm) p 33=2

OpenStudy (saifoo.khan):

it must be equal to 1.

OpenStudy (anonymous):

ok ya that is right

OpenStudy (saifoo.khan):

how can this b 2?

OpenStudy (saifoo.khan):

p(d) = 1 +(d/33) p(0) = 1 +(0/33) p(0) = 1 + 0 p(0) = 1

OpenStudy (anonymous):

k it says the function p(d) =1+d/33 note that p(o)=1 atm p (33)=2 and so on find the pressure at 140 feet

OpenStudy (saifoo.khan):

Yeah, u r correct now.

OpenStudy (saifoo.khan):

it is a Simple prob, As simple as a Cake.

OpenStudy (saifoo.khan):

p(140) = 1 +(140/33) p(140) = 5.24

OpenStudy (saifoo.khan):

got that?

OpenStudy (anonymous):

so the answer would be p(140) =5.24

OpenStudy (anonymous):

wow you are really good can i get you to help me with a few more please

OpenStudy (saifoo.khan):

Ok.

OpenStudy (anonymous):

determine whather the graphs of the two equations are perpendicular y=9x-3 24=3-x

OpenStudy (saifoo.khan):

chk out the 2nd eq.

OpenStudy (saifoo.khan):

there is no y.

OpenStudy (anonymous):

sorry i wrote it wrong 2y=3-x this is right

OpenStudy (anonymous):

so its y=9x-3 2y=3-x are they perpendicular on a graph?

OpenStudy (anonymous):

are you there Saifoo . Khan ?

OpenStudy (saifoo.khan):

hey there.

OpenStudy (saifoo.khan):

??

OpenStudy (saifoo.khan):

Im soooo sorry!

OpenStudy (anonymous):

this one is hard